# The Repeat Offenders: Characterizing and Predicting Extremely Bug-Prone Source Methods

### v1

@@ 6. Threats to Validity
We have identified multiple multiple threats that may impact the validity of our findings.
Construct validity is impacted by our choice of the CodeShovel tool for extracting method histories. Although CodeShovel has been known to show poor performance for very small methods (chowdhury_evidence_2025), it has still outperformed other tools, including FinerGit and IntelliJ / git log (grund_codeshovel_2021). A more recent tool, CodeTracker (jodavi_accurate_2022), which has been shown to outperform CodeShovel, can be used in future work to see how our results hold up.
We used a variety of popular product metrics for evaluating the ExtremelyBuggy methods, however, other metrics could be considered. For example, developer-centric metrics could be used, although we chose not to include developer-centric metrics as they are not very useful at the beginning of a project, and we wanted our findings to be applicable at a projects inception.
@@
To detect the bug-proneness of a method, we used a keyword-based approach, which is known to suffer from false positives (bird2009fair). However, this threat was mitigated by the construction of three different datasets: one with high precision, one with high recall, and one aimed at a balance between precision and recall.
To identify themes in the ExtremelyBuggy methods, we used a manual labeling approach, which in inherently susceptible to human bias and error. To mitigate this threat, we employed the help of two graduate students, with 2 and 10 years of industry experience respectively. After two of the authors labeled 10 methods, these graduate students analyzed the results, and verified the correctness of the assigned labels. The use of two separate authors for manual labeling of methods may also be a threat to validity if they don’t have agreement between the assignment of labels. To curb this threat, the two authors, labeled 10 of the same methods independently, then discussed together and resolved any disagreements that arose. They did this again for another 10 methods, and, after finding almost no disagreements, conducted the rest of the analysis separately.
@@
For the manual analysis of the methods, the high precision dataset was used to reduce the number of false positives during labeling. However, the high precision dataset can only consider a commit to be a bug-fix if there was no tangled changes, that is, if only one method was changed in that commit. This may introduce bias towards certain types of bugs that only affect one method. Future work can be done to investigate ExtremelyBuggy methods bugs that affect multiple methods to validate our results. Additionally, the ExtremelyBuggy methods in the high precision dataset are not evenly distributed, with over half of all ExtremelyBuggy methods coming from just 5 projects. However, upon removing these projects we found that none of the main takeaways from the paper changed, in turn mitigating this threat.
Internal validity is hampered due to our selection of statistical tests such as Kendall’s $\tau$, the Wilcoxon Rank Sum test and Cliffs d. However, all three are well established and used in software engineering research (chowdhury_revisiting_2022; chowdhury_evidence_2025; gil_correlation_2017; inozemtseva2014coverage).
@@
External validity is affected due to our use of only open source, java repositories, and it is unclear how our findings would extend to closed source projects or projects in different languages. We also ensured that the descriptions for our repositories were written in english. This was necessary to ensure bug-fix commits were able to identified using a english language keyword based approach and to ensure we could use commit messages and comments to assist in manual labeling. Future work is needed to find explore the differences in repositories with commit messages written in different languages.
@@
Conclusion validity of this study is influenced by all previous threats to validity. Additionally, our selection of a five-year threshold for age-normalization was chosen arbitrarily to allow for more change history of methods. Consequently, we repeated most of our experiments using a two and eight year threshold, and none of the major results from the first three RQs changed. However, future work is needed to see how the results may change for the manual analysis of ExtremelyBuggy methods using different age-normalization thresholds.
@@ 7. Conclusion
This study provides the first large-scale, method-level investigation into ExtremelyBuggy methods, those repeatedly involved in bug-fix activities, using a dataset of over 1.25 million methods from 98 open-source Java systems. Our empirical analysis demonstrates that although ExtremelyBuggy methods constitute only a small fraction of a project’s methods, they frequently account for a disproportionately large share of bugs. These methods exhibit clear distinguishing characteristics at their inception: they are typically larger, more complex, less readable, and less maintainable than both simple buggy and non-buggy methods. Yet, despite these measurable differences, our evaluation of multiple machine learning models reveals that early prediction of ExtremelyBuggy methods remains highly challenging due to severe class imbalance, cross-project variability, and the fact that many defects arise through later code evolution rather than initial implementation. These findings highlight limitations in existing metric-based prediction approaches and emphasize the need for richer representations of code and project history.
To complement the quantitative analyses, our manual thematic examination of 265 ExtremelyBuggy methods reveals recurring visual, contextual, and defect-related patterns. Visually, these methods often exhibit clear code smells, including confusing control flow, abnormal size growth, poor readability, self-admitted technical debt, and fragile exception-handling structures. Many ExtremelyBuggy methods implement core logic, transform complex data structures, or interface with external resources—roles inherently prone to evolving requirements and error-prone interactions. Common root causes of their repeated failures include missing or incorrect conditional logic, improper exception handling, misuse of variables, and interaction errors stemming from related APIs. These observations carry practical implications: developers may benefit from additional scrutiny and testing of methods that embody these high-risk characteristics, while researchers should explore hybrid models that integrate code embeddings, change-history signals, and contextual semantic information. Overall, this work deepens the understanding of extreme bug-prone methods and establishes a foundation for future research toward more robust, early identification of methods likely to cause recurring bugs.

### v2

@@ 6. Threats to Validity
We have identified multiple threats that may impact the validity of our findings.
@@
Construct validity is impacted by our choice of the CodeShovel tool, which is not 100% accurate while extracting method histories. We used a variety of popular product metrics for evaluating the ExtremelyBuggy methods; however, other metrics could be considered. For example, developer-centric metrics could be used. In this paper, we chose not to include developer-centric metrics as they are not very useful at the beginning of a project, and we wanted our findings to be applicable at a project’s inception. To detect the bug-proneness of a method, we used a keyword-based approach, which is known to suffer from false positives (Bird et al., 2009). However, this threat was mitigated, to some extent, by the construction of three different datasets: one with high precision, one with high recall, and one aimed at a balance between precision and recall.
@@
For the manual analysis of the methods, the HighPrecision dataset was used to reduce the number of false positives during labeling. However, the HighPrecision dataset can only consider a commit to be a bug-fix if there were no tangled changes, that is, if only one method was changed in that commit. This may introduce bias towards certain types of bugs that only affect one method.
Internal validity is hampered due to our selection of statistical tests such as Kendall’s $\tau$, the Wilcoxon Rank Sum test, and Cliff’s d. However, these tests are well established and commonly used in software engineering research (Chowdhury et al., 2022b; Gil and Lalouche, 2017; Inozemtseva and Holmes, 2014). To identify themes in the ExtremelyBuggy methods, we used a manual labeling approach, which is inherently susceptible to human bias and error. To mitigate this threat, we compared the labeling of the first two authors and also received help from the third and fourth authors with 2+ and 10+ years of industry experience, respectively.
@@
External validity is affected by our use of only open-source Java repositories, and it is unclear how our findings would extend to closed-source projects or projects in other languages. Also, for our analysis, we had to ensure that the descriptions (e.g., commit messages) for our repositories were written in English, and discarded other repositories with non-English descriptions.
@@
Conclusion validity of this study is influenced by all previous threats to validity. Additionally, our selection of a five-year threshold for age-normalization can impact the conclusion we made. To mitigate this threat, we repeated most of our experiments using a two and eight-year threshold, and none of the major results from the first three RQs changed. However, future work is needed to see how the results may change for the manual analysis of ExtremelyBuggy methods using different age-normalization thresholds.
@@ 7. Conclusion
This study provides the first large-scale, method-level investigation into ExtremelyBuggy methods using a dataset of over 1.25 million methods from 98 open-source Java systems. Our empirical analysis demonstrates that although ExtremelyBuggy methods constitute only a small fraction of a project’s methods, they frequently account for a disproportionately large share of bugs. These methods exhibit clear distinguishing characteristics at their inception: they are typically larger, more complex, less readable, and less maintainable than both simple buggy and non-buggy methods. Yet, despite these measurable differences, our evaluation of multiple machine learning models reveals that early prediction of ExtremelyBuggy methods remains highly challenging due to tangled code changes, severe class imbalance, cross-project variability, and the fact that many defects arise through later code evolution rather than initial implementation. To complement the quantitative analyses, our manual thematic examination of 265 ExtremelyBuggy methods reveals recurring visual, contextual, and defect-related patterns. Visually, these methods often exhibit clear code smells, including confusing control flow, abnormal size growth, poor readability, self-admitted technical debt, and fragile exception-handling structures.
These findings emphasize the need for richer representations of code and project history and the necessity of more research on how to deal with class imbalance and tangled changes. Overall, our work deepens the understanding of extreme bug-prone methods and establishes a foundation for future research toward more robust, early identification of methods likely to cause recurring bugs.
Acknowledgements.
This research project was supported by an NSERC Discovery Grant, Canada.